# Yiqi Lin1* Conghui He1*† Alex Jinpeng Wang2* Bin Wang1* Weijia Li3 Mike Zheng Shou2

### 21 Dec 2023 v1

@@ 2.2 Studying of CLIP Behaviors
Despite the strong zero-shot and transferable performance of CLIP, the perpetuating biases [8, 30, 1, 34, 14] in CLIP are still not well investigated due to its large-scale noisy training data. Much research [32, 18, 27] focuses on revealing or enhancing the downstream performance with discovered bias in CLIP. For example, colorful masks [32] or red circles [27] applied to images can improve the zero-shot performance on visual grounding or keypoint localization tasks. In studying visual text content bias,  [8] shows the multimodal neurons of CLIP not only respond to visual content and the visual text embedded in the image. Another work [14] shows that image recognition in CLIP can be strongly dominated by the visual text embedded in the image. To disentangle such bias, [18] attempts to separate the text spotting representation in pre-trained CLIP by training representation projection. Meanwhile, LoGoPrompt [26] enhances the classification performance by utilizing the visual text content as auxiliary prompts as input. Also, CLIPPO [29] shows that directly aligning the image and synthetic images embedded with the captions can perform similarly to CLIP without a text-specific encoder. Differently, this paper makes the first attempt to reveal the source of text spotting capacity in CLIP is the data bias.
@@ 3 Terminology
Algorithm 1 Pseudocode of Detecting Co-Emb. Text (Rate)
# caption: captions from LAION-2B dataset.
# ocr_text: text spotted by OCR model.
cap_words = set(caption.split())
ocr_words = set(ocr_text.split())
co_emb_text = intersection(cap_words, ocr_words)
co_emb_text_rate = len(co_emb_text) / len(cap_words)
@@ 5.1 Ablation of Embedded Text Removal
6 Training CLIP on Emb. Text Curated Data
6.1 Experiment Setups
@@ 6.2 Ablation Study
Number of Total Images 323746044
Number of Images w/ Emb. Text 162379774
Image w/ Emb. Text Rate 50.15%
@@
– (in Images w/ Emb. Text) 15.68%
Fuzzy Co-Emb. Text Rate (in Total) 5.74%
– (in Images w/ Emb. Text) 31.26%

### 28 Dec 2023 v2

@@ 2.2 Studying of CLIP Behaviors
Despite the strong zero-shot and transferable performance of CLIP, the perpetuating biases [9, 33, 1, 38, 15] in CLIP are still not well investigated due to its large-scale noisy training data. Much research [36, 20, 30, 35] focuses on revealing or enhancing the downstream performance with discovered bias in CLIP. For example, colorful masks [36] or red circles [30] applied to images can improve the zero-shot performance on visual grounding or keypoint localization tasks. In studying visual text content bias,  [9] shows the multimodal neurons of CLIP not only respond to visual content and the visual text embedded in the image. Another work [15] shows that image recognition in CLIP can be strongly dominated by the visual text embedded in the image. To disentangle such bias, [20] attempts to separate the text spotting representation in pre-trained CLIP by training representation projection. Meanwhile, LoGoPrompt [29] enhances the classification performance by utilizing the visual text content as auxiliary prompts as input. Also, CLIPPO [32] shows that directly aligning the image and synthetic images embedded with the captions can perform similarly to CLIP without a text-specific encoder.
@@
2.3 Data Curation with Text Removal
Due to the successful practice of data curation in LAION datasets [27, 28] on scaling up the image-text datasets, searching advanced selection strategy to improve the data quality from common crawl data pool gains a growing interest [8]. Recently, several works [24, 19, 4] suggest that introducing text-related filtering methods improves the pre-training dataset quality. In DiHT [24], the data curation steps include filtering out the image-text pairs with high OCR confidence and matching text ratio. Moreover, [19, 4] mainly focus on studying the importance of filtering out the text-dominate images utilizing OCR models to improve pre-training dataset quality. Maini et al. [19] also draw the observation that 40% of LAION’s image text is highly correlated with the caption, but only performing a small pilot study on 500 samples with manual judgment. Differently, this paper makes the first attempt to reveal the source of text spotting capacity in CLIP is the data bias and the consequences of such bias in existing commonly used datasets.
Algorithm 1 Pseudocode of Detecting Co-Emb. Text (Rate)
# caption: captions from LAION-2B dataset.
# ocr_text: text spotted by OCR model.
cap_words = set(caption.split())
ocr_words = set(ocr_text.split())
co_emb_text = intersection(cap_words, ocr_words)
co_emb_text_rate = len(co_emb_text) / len(cap_words)
@@ 5.1 Ablation of Embedded Text Removal
6 Training CLIP on Emb. Text Curated Data
6.1 Experiment Setups
@@ 6.2 Ablation Study
Number of Total Images 516166764
Number of Images w/ Emb. Text 259101894
Image w/ Emb. Text Rate 50.20%
@@
– (in Images w/ Emb. Text) 15.70%
Fuzzy Co-Emb. Text Rate (in Total) 5.75%
– (in Images w/ Emb. Text) 31.28%